# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ 5.3.3 Performance Analysis Across Scan patterns (Pooled Over All Teams)
The trends observed across figures above consistently demonstrate that both global and local error metrics increase with scan length, regardless of scan patterns (e.g., protocol, arm, or orientation). This suggests a strong dependence of metric magnitude on scan length. To validate this observation quantitatively, the correlation between scan length and each metric is assessed using the Pearson correlation coefficient ($r$). For most metrics, $r$ values range from 0.3 to 0.5, indicating moderate positive correlations between scan length and error magnitude. Notably, the four subject-level metrics exhibit stronger correlations, with $r$ values between 0.58 and 0.78, suggesting a substantial association between scan length and performance degradation at the subject level. In all cases, the correlations are statistically significant ($p<0.05$).
@@ 5.3.4 Qualitative Results
Fig. 22 presents the qualitative results of the participating teams, illustrating scans corresponding to the best, worst, and median performance based on GPE (subfigures a-c) and LPE (subfigures d-f) metric errors, while accounting for variability across scan patterns. To ensure representative coverage and avoid redundant selection of the same scan pattern, the scan with the next-lowest (or next-highest / next-median) error was selected when appropriate. To show the quantitative difference, the corresponding numerical metric values are provided. Fig. 22 illustrates that accurate local predictions do not necessarily ensure accurate global reconstructions. Additionally, the error magnitude varies across different scans.
@@
Figure 17: Distribution of error metrics (GPE, GLE, LEP, LLE) for individual scans across all methods, categorised by probe orientation (parallel and perpendicular) and ordered by increasing scan length within each group.
@@ 5.3.5 Further Analysis
Across all evaluated methods, a primary limitation is the sensitivity to scan length, as reflected by the correlation between scan length and error magnitude. This suggests that most approaches struggle to maintain lower global error over long sequences, likely due to cumulative drift or limited global context modeling. Furthermore, performance degradation is particularly prominent in geometrically complex trajectories (e.g., S shape scans), indicating an overall lack of robustness to scanning path variability.
While all teams share these global limitations, individual methods demonstrate distinct advantages. Some teams excel in local consistency, achieving low LPE and LLE across a wide range of scans, indicating effective frame-to-frame alignment strategies. Others achieve stronger performance on long scans, suggesting better global modeling. However, no team consistently outperform others across all settings. This lack of generalisation highlights the current trade-offs between local precision and global robustness.
Notably, the positive impact of ensemble learning and pretraining highlights the value of integrating prior visual knowledge, which could be considered when clinically deployed. Taken together, these results suggest that future development should prioritise hybrid approaches that combine local precision, temporal consistency, and robust drift correction. Additionally, performance across scan patterns emphasises the need for improving adaptability in clinical scanning environments.
@@ 6 Discussions
While the current Challenge setup offers a comprehensive evaluation of freehand 3D ultrasound reconstruction methods, several limitations remain and highlight opportunities for improvement in future iterations.
Reducing barriers to participation. A key priority for future editions of the TUS-REC Challenge is to lower the technical entry barrier, thereby enabling broader participation from research groups across computer vision, robotics, and medical imaging domains, including those without prior experience in freehand ultrasound reconstruction. Currently, the prerequisite knowledge of ultrasound imaging principles, spatial calibration procedures, and coordinate systems may discourage otherwise capable participants. By providing detailed documentations, the Challenge can broaden accessibility and encourage participation from a more diverse range of research communities. Notably, with existing effort provided by this Challenge paper, this barrier has already been significantly lowered, laying a strong foundation for continued growth and engagement.
@@
Score weighting and ranking methodologies. In TUS-REC2024 Challenge, a min-max normalisation strategy is applied at the scan level, scaling performance scores into a fixed range of $[0,1]$. While this approach helps reduce the influence of extreme values, it introduces several limitations. For example, it can over-amplify marginal differences in performance when overall variability is low, leading to exaggerated score separation between similarly performing methods. Additionally, the presence of a poorly performing team can artificially boost the normalised scores of others with only slightly better performance. These effects may distort the overall fairness and interpretability of rankings. It is important to acknowledge that all normalisation strategies inherently involve trade-offs, and no single method is universally optimal across all evaluation contexts. Min-max normalisation can suppress the impact of outliers but may exaggerate differences when overall performance is similar. In contrast, z-score normalisation mitigates sensitivity to extreme values but may obscure meaningful differences when the distribution of scores is non-Gaussian. Ultimately, each method emphasises different aspects of relative performance, and the choice of normalisation inevitably shapes the interpretation of results. While careful selection and justification of normalisation methods can improve transparency, there is no definitive solution that fully eliminates bias or distortion in score scaling. Therefore, normalisation should be viewed as a practical necessity with acknowledged limitations, rather than a universally fair metric transformation.
Risk of data leakage. As the Challenge allows multiple submission attempts and uses a fixed test set, there is a risk that participating methods may become inadvertently overfitted to the test data over time. This is especially relevant when teams refine their models based on feedback from repeated evaluations, potentially optimising for the specifics of the test distribution rather than generalisable performance. Such overfitting can undermine the fairness and validity of the ranking results. To mitigate this issue, future editions of the TUS-REC Challenge should consider introducing additional unseen test data in later evaluation phases. This could include a hold-out set only revealed after the final submission deadline or a progressive test set release strategy. Incorporating fresh data would better evaluate the generalisation ability of submitted methods and reduce the likelihood of overfitting to a static benchmark. Moreover, it would more closely reflect real-world deployment conditions, where models must perform reliably on previously unseen patients and scanning conditions.
@@ 7 Conclusion
Trackerless 3D freehand ultrasound reconstruction represents a critical advancement in enabling cost-effective, portable, and workflow-friendly 3D imaging solutions for diverse clinical settings. By eliminating the need for external tracking hardware, these methods promise improved accessibility in point-of-care diagnostics and interventional guidance. However, this paradigm shift also introduces new algorithmic challenges in accurate motion estimation under unconstrained probe motion.
TUS-REC2024 Challenge represents a major step forward in benchmarking the current state of the art for trackerless 3D freehand ultrasound reconstruction. With the largest publicly available dataset for this task and participation from leading international research teams, it has successfully enabled a comparative evaluation of modern methods under a standardised test framework. The dataset, containing over two thousand of scans across multiple subjects, provides an invaluable resource for the community and will continue to support method development and reproducibility beyond the Challenge.
The submitted methods reflect a rich diversity of algorithmic strategies, including spatial and temporal modeling, data augmentation, and fusion architectures. These contributions offer a strong foundation for future research. Despite notable advancements, the task is not yet solved to a clinically satisfactory degree. Sensitivity to scan length and scanning protocol reveals that generalisation remains a challenge, and further work is needed to bridge the gap between experimental performance and clinical deployment.
@@
The training and validation datasets used in TUS-REC2024 Challenge are publicly available at the following repositories: https://doi.org/10.5281/zenodo.11178508, https://doi.org/10.5281/zenodo.11180794, https://doi.org/10.5281/zenodo.11355499, and https://doi.org/10.5281/zenodo.12752245. The baseline model code has been released and can be accessed via GitHub at https://github.com/QiLi111/tus-rec-challenge_baseline. Participants are encouraged to release their code voluntarily. All publicly released code related to the Challenge will be listed on the official Challenge website: https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/.
@@ Acknowledgement
This work was supported by the EPSRC [EP/T029404/1], a Royal Academy of Engineering/Medtronic Research Chair [RCSRF1819\7\734] (TV), Wellcome/EPSRC Centre for Interventional and Surgical Sciences [203145Z/16/Z], and the International Alliance for Cancer Early Detection, an alliance between Cancer Research UK [C28070/A30912; C73666/A31378], Canary Center at Stanford University, the University of Cambridge, OHSU Knight Cancer Institute, University College London and the University of Manchester. TV is co-founder and shareholder of Hypervision Surgical. Qi Li was supported by the University College London Overseas and Graduate Research Scholarships. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 5.3.3 Aggregated Performance over All Teams and Influence of Scan Length
The trends observed across figures above consistently demonstrate that both global and local error metrics increase with scan length, regardless of scan patterns (e.g., protocol, orientation or direction). This suggests a strong dependence of metric magnitude on scan length. To validate this observation quantitatively, the correlation between scan length and each metric is assessed using the Pearson correlation coefficient ($r$). For most metrics, $r$ values range from 0.3 to 0.5, indicating moderate positive correlations between scan length and error magnitude. Notably, the four subject-level metrics exhibit stronger correlations, with $r$ values between 0.58 and 0.78, suggesting the association between scan length and performance degradation at the subject level. In all cases, the correlations are statistically significant ($p<0.05$).
@@ 5.3.4 Qualitative Results
Fig. 23 plots the predicted trajectories for each participating teams, selecting the scans with the best, worst, and median performance based on average GPE (subfigures a-c) and average LPE (subfigures d-f) of all teams, respectively. To include as diverse scan patterns as possible so that representativeness can be guaranteed, the scan with the next-lowest (or next-highest / next-median) error was selected until the three cases selected based on each metric have different types of scan shapes. To show the quantitative difference, the corresponding GPE and LPE values are also provided. Fig. 23 illustrates that accurate local predictions do not necessarily ensure accurate global reconstructions. In addition, the error magnitude varies across different scans.
@@
Figure 18: Distribution of error metrics (GPE, GLE, LPE, LLE) for individual scans across all methods, categorised by probe orientation (parallel and perpendicular) and ordered by increasing scan length within each group.
@@ 5.3.5 Further Analysis
Across all evaluated methods, a primary limitation is the sensitivity to scan length, as reflected by the correlation between scan length and error magnitude. This suggests that most approaches struggle to maintain lower global error over long sequences, likely due to lack of long range correlation. Furthermore, performance degradation is particularly prominent in geometrically complex trajectories (e.g., S shape scans), indicating an overall lack of robustness to scanning path variability.
While all teams share these global limitations, individual methods demonstrate distinct advantages. Some teams excel in local consistency, achieving low LPE and LLE across a wide range of scans, indicating effective frame-to-frame alignment strategies. Others achieve stronger performance on long scans, suggesting better global modeling. However, no team consistently outperform others on all metrics. This inconsistency across metrics highlights the current trade-offs between local precision and global robustness.
Notably, the performance gain brought by ensemble learning and pretraining highlights the value of integrating uncertainty calibration and prior visual knowledge, which could be considered when clinically deployed. Taken together, these results suggest that future development should prioritise hybrid approaches that combine local precision, temporal consistency, and robust drift correction. Additionally, performance difference across scan patterns emphasises the need for improving adaptability in clinical scanning environments.
@@ 6 Discussions
While the current Challenge setup offers a comprehensive evaluation of trackerless freehand 3D ultrasound reconstruction methods, several limitations remain, which also highlight opportunities for improvement in future iterations.
Reducing barriers to participation. A key priority for future editions of the TUS-REC Challenge is to lower the technical entry barrier, thereby enabling broader participation from research groups across computer vision, robotics, and medical imaging domains, including those without prior experience in freehand ultrasound reconstruction. Currently, the prerequisite knowledge of ultrasound imaging principles, spatial calibration procedures, and coordinate systems may discourage otherwise capable participants. By providing more detailed documentations, the Challenge can broaden accessibility and encourage participation from a more diverse range of research communities. Notably, with existing effort provided by this Challenge paper, this barrier has already been significantly lowered, laying a strong foundation for continued growth and engagement.
@@
Score weighting and ranking methodologies. In TUS-REC2024 Challenge, a min-max normalisation strategy is applied at the scan level, scaling performance scores into a fixed range of $[0,1]$. While this approach helps reduce the influence of extreme values, it introduces several limitations. For example, it can over-amplify marginal differences in performance when overall variability is low, leading to exaggerated score separation between similarly performing methods. Additionally, the presence of a poorly performing team can artificially boost the normalised scores of the other teams. These effects may distort the overall fairness and interpretability of rankings. It is important to acknowledge that all normalisation strategies inherently involve trade-offs, and no single method is universally optimal across all evaluation contexts. Min-max normalisation can suppress the impact of outliers but may exaggerate differences when overall performance is similar. In contrast, z-score normalisation avoids exaggerated differences but cannot suppress the impact of outliers as min-max normalisation does. Ultimately, each method emphasises different aspects of relative performance, and the choice of normalisation inevitably shapes the interpretation of results. While careful selection and justification of normalisation methods can improve transparency, there is no definitive solution that fully eliminates bias or distortion in score scaling. Therefore, normalisation should be viewed as an empirical approach with acknowledged limitations.
Risk of data leakage. As the Challenge allows multiple submission attempts and uses a fixed test set, there is a risk that participating methods may become inadvertently overfitted to the test data over time. This is especially relevant when teams refine their models based on feedback from repeated evaluations, potentially optimising for the specific test distribution rather than generalisable performance. Such overfitting can undermine the fairness and validity of the ranking results. To mitigate this issue, future editions of the TUS-REC Challenge should consider introducing additional unseen test data in later evaluation phases. This could include a hold-out set only revealed after the final submission deadline or a progressive test set release strategy. Incorporating fresh data would better evaluate the generalisation ability of submitted methods and reduce the likelihood of overfitting to a static benchmark. Moreover, it would more closely reflect real-world deployment conditions, where models must perform reliably on previously unseen patients and scanning conditions.
@@ 7 Conclusion
Trackerless 3D freehand ultrasound reconstruction represents a critical advancement in enabling cost-effective, portable, and workflow-friendly 3D imaging solutions for diverse clinical settings. By eliminating the need for external tracking hardware, these methods promise improved accessibility in point-of-care diagnostics and interventional guidance. However, it also introduces new algorithmic challenges in accurate motion estimation under unconstrained probe motion.
TUS-REC2024 Challenge represents a major step forward in benchmarking the current state-of-the-art for trackerless 3D freehand ultrasound reconstruction. With the large publicly available dataset for this task and participation from leading international research teams, it has successfully enabled a comparative evaluation of modern methods under a standardised test framework. The dataset, containing over two thousand scans across multiple subjects, provides an invaluable resource for the community and will continue to support method development and reproducibility beyond the Challenge.
The submitted methods reflect a rich diversity of algorithmic strategies, including spatial and temporal modeling, data augmentation, and ensemble learning. These contributions offer a strong foundation for future research. Despite notable advancements, this task is not yet solved to a clinically satisfactory degree. Sensitivity to scan length and scan pattern reveals that generalisation remains a challenge, and further work is needed to bridge the gap between experimental performance and clinical deployment.
@@
The training and validation datasets used in TUS-REC2024 Challenge are publicly available at the following repositories: https://doi.org/10.5281/zenodo.11178508, https://doi.org/10.5281/zenodo.11180794, https://doi.org/10.5281/zenodo.11355499, and https://doi.org/10.5281/zenodo.12752245. The baseline model code has been released and can be accessed via GitHub at https://github.com/QiLi111/tus-rec-challenge_baseline. Participants are encouraged to release their code voluntarily. All publicly released codes related to the Challenge are listed on the official Challenge website: https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/leaderboard.html.
@@ Acknowledgement
This work was supported by the EPSRC [EP/T029404/1], a Royal Academy of Engineering/Medtronic Research Chair [RCSRF1819\7\734] (TV), Wellcome/EPSRC Centre for Interventional and Surgical Sciences [203145Z/16/Z], and the International Alliance for Cancer Early Detection, an alliance between Cancer Research UK [C28070/A30912; C73666/A31378; EDDAPA-2024/100014], Canary Center at Stanford University, the University of Cambridge, OHSU Knight Cancer Institute, University College London and the University of Manchester. TV is co-founder and shareholder of Hypervision Surgical. Qi Li was supported by the University College London Overseas and Graduate Research Scholarships. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.